OpenStudy (anonymous):

\frac{g}{-5} +3 = -13

jimthompson5910 (jim_thompson5910):

your first step is to subtract 3 from both sides

jimthompson5910 (jim_thompson5910):

\[\large \frac{g}{-5} +3 = -13\] \[\large \frac{g}{-5} +3-3 = -13-3\] \[\large \frac{g}{-5} +0 = -16\] \[\large \frac{g}{-5} = -16\] what's next?

OpenStudy (anonymous):

what would G equal?

jimthompson5910 (jim_thompson5910):

to find out, you need to multiply both sides by -5

jimthompson5910 (jim_thompson5910):

because you want to undo that division of -5

OpenStudy (anonymous):

okay hold on

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

is it 80?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

THANK YOU!

jimthompson5910 (jim_thompson5910):

you're welcome